### v2

# Frontier AI’s Impact on the Cybersecurity Landscape

Abstract
As frontier AI advances rapidly, understanding its impact on cybersecurity and inherent risks is essential to ensuring safe AI evolution (e.g., guiding risk mitigation and informing policymakers). While some studies review AI applications in cybersecurity, none of them comprehensively discuss AI’s future impacts or provide concrete recommendations for navigating its safe and secure usage. This paper presents an in-depth analysis of frontier AI’s impact on cybersecurity and establishes a systematic framework for risk assessment and mitigation. To this end, we first define and categorize the marginal risks of frontier AI in cybersecurity and then systemically analyze the current and future impacts of frontier AI in cybersecurity, qualitatively and quantitatively. We also discuss why frontier AI likely benefits attackers more than defenders in the short term from equivalence classes, asymmetry, and economic impact. Next, we explore frontier AI’s impact on future software system development, including enabling complex hybrid systems while introducing new risks. Based on our findings, we provide security recommendations, including constructing fine-grained benchmarks for risk assessment, designing AI agents for defenses, building security mechanisms and provable defenses for hybrid systems, enhancing pre-deployment security testing and transparency, and strengthening defenses for users. Finally, we present long-term research questions essential for understanding AI’s future impacts and unleashing its defensive capabilities.
11footnotetext: These authors contributed equally to this work.
1 Introduction
Frontier artificial intelligence (AI), including foundation models and emerging AI agents, has demonstrated remarkable success across domains such as text generation [2], code generation and debugging [113, 236], scientific discovery [234, 149], and healthcare [44]. Recent works [115, 85, 312] further explore AI agents to accomplish complex tasks beyond content generation, such as automatically resolving software issues [312]. While “frontier AI” has been used to refer to foundation models only [11], we use it to represent broader advanced AI systems, including foundation models like large language models (LLMs) [2, 276, 272, 15] and multi-modal models [293], as well as AI agents [85]. [^1] We use both AI agents and hybrid systems to refer to systems that combine foundation models with non-ML symbolic software components, where foundation models interact with other components.
Motivated by their strong capabilities, the security community has also begun utilizing frontier AI, such as using LLM to facilitate fuzzing [196] and generate patching [332]. Some works further explored building foundation models for security data, such as networking traffic, binary code, and blockchain transactions, achieving better performance than traditional ML in threat detection [117, 98, 320, 220]. Recently, AI-driven cyber reasoning systems from the DARPA AIxCC competition [71] found and patched real-world zero-day vulnerabilities [119]. While frontier AI offers great promise, it also introduces significant risks. First, attackers can leverage LLMs to synthesize malicious code [151, 286] or to launch attacks against systems and humans [90, 271]. Second, AI models themselves have serious security concerns [287, 26], including backdoors [205], jailbreaking [52, 53], prompt injection [181], and data leakage [41].
Figure 1: Performance evolvement of frontier AI models on coding, math, and science benchmarks over time. The coding-related tasks are emphasized in bold text.
Given these successes and concerns, a key question arises: Will frontier AI benefit attackers or defenders more, and how might it reshape the cybersecurity landscape? While recent work surveys LLMs’ applications in cybersecurity [93, 124, 102, 318, 239, 331, 337, 136, 253], they do not provide a comprehensive analysis of how frontier AI will impact the cybersecurity landscape. These works also do not analyze potential risks introduced by frontier AI or provide actionable recommendations on how the security community should react to such risks.
With the rapid AI development, it is critical to analyze frontier AI’s security impact and provide recommendations [1, 249]. Figure 1 shows frontier AI models’ capabilities improving at an unprecedented pace, particularly in code reasoning, generation, and patching. The o3 model—launched just two months after o1—significantly outperformed o1 on Codeforces (2727 vs. 1891) and the SWE-bench Verified benchmark (71.7% vs. 48.9%) [278]. This trend implies that AI could become increasingly leverageable by attackers than previously expected.
In this paper, we address the gap by presenting a comprehensive study of frontier AI’s current and future impact on cybersecurity, as well as a call to action for system builders, defenders, and AI developers. We begin by defining and analyzing the marginal risks [153] of frontier AI to cybersecurity and provide a detailed risk categorization based on the standard cyber kill chain [86]. Using this categorization, we summarize frontier AI’s current use and capability in attacks and defenses and discuss its potential future impacts, quantitatively and qualitatively. Next, we analyze whether frontier AI will benefit attackers or defenders more from the perspectives of equivalence classes, asymmetries between attackers and defenders, and economic impacts. We then explore how frontier AI will reshape future system design. Based on these projections, we offer actionable recommendations for leveraging frontier AI to build more secure systems while assessing and preparing for its potential risks. Finally, we present a set of open research questions about frontier AI and security.
Findings.
  • Our systematic review shows that most AI-driven attacks on software systems are demonstrated in research settings and are not yet widely applied in practice. Conversely, many real-world AI-driven attacks on humans (e.g., phishing, social engineering, deepfakes) are already occurring on a large scale.
  • Frontier AI advances automation in defense, particularly in threat detection and Security Operations Center (SOC) operations, but has not been widely used in vulnerability triage and remediation. Overall, its influence on defense is less than on offense.
  • We argue that frontier AI will likely benefit attackers more than defenders in the short term. However, with a coordinated effort—through enhanced risk assessment, smarter defense integration, secure-by-design system development—the balance could eventually tip in favor of defenders.
  • We expect more hybrid systems combining AI with traditional symbolic components, which will introduce new security challenges.
Call to action.
  • We recommend comprehensive risk assessments and continuous monitoring of frontier AI capabilities and resulting cybersecurity risks. This involves building comprehensive, fine-grained benchmarks, dynamically updating them, and continuously conducting marginal risk assessments.
  • We recommend strategically leveraging frontier AI to strengthen every phase of the cybersecurity defense lifecycle. To this end, we can build defenses combining AI with existing program analysis-based approaches. We can also explore hybrid systems with multiple AI agents to enhance defenses.
  • We recommend more efforts to foster better security for hybrid systems, whether through secure-by-design approaches or by creating specialized proactive/reactive defenses.
  • We recommend AI developers conduct more comprehensive pre-deployment security testing and enforce greater transparency in AI development. Moreover, giving defenders early and differential access can help shift the balance in favor of the defenders over the attackers.
  • To mitigate human-targeted attacks, we call for better education about AI-enabled attacks and security best practices, while advancing research on defenses against sophisticated AI-powered threats targeting humans.
These recommendations aim to guide the community to understand and mitigate frontier AI’s marginal risks, turning it into a powerful tool for strengthening cybersecurity.
2 Methodology
2.1 Marginal Risk Assessment Framework
Based on the concept of marginal risk [153], which refers to new and unique risks introduced by emerging technologies, we define the marginal risks of frontier AI in cybersecurity as novel cyber threats either introduced or amplified by frontier AI through intentional misuse beyond traditional attack methods.
First, we study how frontier AI amplifies existing attacks against systems and users. For attacks against systems, we follow the MITRE ATT&CK framework [201] and cyber kill chain [190] (Figure 6) to categorize existing attacks into seven steps. We follow the cyber kill chain to name each step and merge the corresponding steps in MITRE ATT&CK into the sub-steps (Figure 2). Similarly, we classify attacks against humans based on their goals and methods (Figure 3). Based on this categorization, we conduct a detailed analysis of the marginal risks of frontier AI to each attack step through qualitative and quantitative analysis.
Second, we follow a similar methodology to analyze frontier AI’s impacts on defenses. We follow the convention to categorize defenses for systems and humans (Figure 4) and analyze each defense step through qualitative and quantitative analysis. Third, with a comprehensive understanding of frontier AI’s impact on attacks and defenses, we predict whether it will benefit attacks or defenses more. Finally, we discuss the marginal risks in hybrid systems (defined in Section 1) and potential mitigations.
We acknowledge that covering all risks and attack vectors is challenging and thus focus on critical attacks that have been impacted by AI. For example, we do not discuss attack vectors in cryptography, such as discovering vulnerabilities in cryptographic protocols. These attack vectors are currently beyond AI’s capabilities, though AI may be useful to such attacks later.
2.2 Qualitative & Quantitative Risk Analysis
For each attack and defense steps, our qualitative analysis exhaustively searches and systematically catalogs AI-facilitated techniques and instances in research papers and real-world deployment from the following four-tier impact: (1) no demonstrated effect, where we find no AI-related impact for this step; (2) demonstration in research papers, where we find only research papers but no real-world reports; (3) demonstration in real-world attacks, where we find research papers and a few early-stage real-world demonstrations; and (4) large-scale real-world deployments, where AI has been widely deployed in the real world.
Beyond qualitative analysis, we further study quantifying frontier AI’s impact in attacks and defenses. To this end, we collect existing benchmarks on AI-facilitated attacks and defenses and summarize their evaluation results. We also discuss the limitations of current benchmarks and how to comprehensively and accurately quantify frontier AI’s impact in the future. Building upon our qualitative and quantitative analysis, we explore the future impacts of frontier AI in cybersecurity and provide concrete recommendations for mitigating potential negative impacts. Table I presents a summary of the status quo.
TABLE I: Key summaries of the current status quo for each attack and defense step. The colors of the first column represent different levels of impact for the corresponding attack/defense step: Green–no demonstrated effect, Yellow–demonstrated in research papers, Orange–demonstrated in the real world, and Red–the large-scale real world.
  Attack steps Summary
Step 1: Reconnaissance (Demonstrated in the real world) Qualitative: Both research and real-world attacks leverage LLMs in reconnaissance, including active scanning [270], victim information gathering [73, 200], and open-source database search [231].
Quantitative: Current benchmarks rely primarily on text-based assessments [274, 184, 286], while practical evaluations like AutoPenBench [101] show AI agents’ potential in reconnaissance but still limited.
Step 2: Weaponization (Demonstrated in the real world) Qualitative: Research papers show LLMs can generate functional malware with high evasion rates [213] and AI agents aid in vulnerability identification and exploitation [186, 174, 90, 91], although large-scale real-world attacks remain limited to simpler attacks like DDoS [23, 281] and SQL injection [200, 119].
Quantitative: RedCode benchmark [111] reveals LLMs’ limitations in generating functional malware. Vulnerability exploitation benchmarks [286, 316, 251, 329] show Claude 3.5 Sonnet achieving cybersecurity practitioner-level skills but falling short of expert capabilities.
Qualitative: Initial access and exploitation (installation) are demonstrated in real world [96, 208, 142] while persistence remains at research level [245].
Step 3-5: Delivery & exploitation & installation (Demonstrated in the real world) Quantitative: Benchmark coverage and quality for these attack steps are very limited.
Step 6: Command and control (Demonstrated in the real world) Qualitative: Most sub-categories have been demonstrated only in research papers, e.g., privilege escalation through exploit chain generation [68] and command and control through automated domain generation [245, 13], while credential access is used in real world [244, 148, 211].
Quantitative: Benchmark coverage and quality for this attack step are very limited.
Step 7: Action on objectives (Large-scale real world impact) Real-world AI-enhanced attacks increase across various systems (Web, mobile, cloud) [23, 137, 63], with malicious purposes including malware deployment, business logic abuse, and credential theft, causing significant financial losses and data breaches [142, 191, 63].
Attacks against humans (Large-scale real world impact) Frontier AI escalates attacks against humans, with studies showing increases in social engineering (135%) and voice phishing (260%) since ChatGPT’s adoption [67, 262, 131]. AI is misused for identity theft [192], deepfake-based fraud [49, 189], child exploitation [266], and psychological manipulation [238].
  Defense steps Summary
Step 1: Proactive testing (Demonstrated in the real world) Qualitative: Research explores LLMs for proactive cybersecurity testing, including automated penetration testing [155, 73, 120] and vulnerability detection through code foundation models [80, 186] and AI-enhanced fuzzing [139, 310]. Real-world demonstrations exist but lack evidence of large-scale adoption.
Quantitative: AI penetration testing benchmarks remain nascent, with one end-to-end benchmark AutoPenBench [101]. Vulnerability detection benchmarks [80, 247, 199] show SOTA models achieve limited accuracy, but these benchmarks face challenges in data quality, limited code context, and insufficient test diversity.
Step 2: Attack detection (Demonstrated in the real world) Qualitative: Frontier AI addresses key limitations of traditional AI methods by eliminating manual feature engineering, reducing dependence on labeled datasets [282, 77], and improving generalizability to out-of-distribution data [117, 7]. Real-world applications exist in malware and network intrusion detection.
Quantitative: Many benchmarks exist for network intrusion and malware detection [252, 19, 315] but have data and label quality limitations. Foundation model-driven detection shows high accuracy [117, 178], calling for more challenging benchmarks.
Step 3: Triage & forensic (Demonstrated in research papers) Qualitative: Frontier AI usages are at an early stage, including LLMs for vulnerability analysis, improving symbolic execution with PoC generation [292, 126] and developing AI agents for root cause analysis [235]. Pre-trained foundation models demonstrate superior performance in binary analysis tasks [220, 305].
Quantitative: Recent benchmarks like CRUXEval [108] show promising results for PoC generation with GPT-4o achieving 75% pass@1 success rate, while no public benchmark exists for comprehensive AI-driven reverse engineering evaluation.
Step 4-5: Remediation development & deployment (No demonstrated effect) Qualitative: Research demonstrates frontier AI can automatically generate security vulnerability patches [177, 159, 3, 34], though real-world applications like the SQLite3 Off-by-One bug fix [119] remain limited and no specific work exists on AI-assisted remediation deployment.
Quantitative: SOTA systems using Claude 3.5 Sonnet resolve about 50% of issues in SWE-bench-verified [147]. However, most SWE-bench issues are non-security bugs, revealing limitations in evaluating frontier AI’s defense capabilities.
Defense for humans (Demonstrated in the real world) Frontier AI shows promise for enhancing defenses against human-targeted attacks, including social bot detection [94], fraud detection [326, 78], deepfake detection [229], and misinformation detection [14, 79, 180]. However, defensive techniques struggle to keep pace with sophisticated attacks [6], highlighting the need for research on adversarial dynamics between defensive and attack AI systems.
3 Frontier AI’s Impact on Cyberattacks
3.1 Attacks Targeting Systems
3.1.1 Qualitative impact analysis
Figure 2 shows frontier AI’s qualitative impact on attack steps. Most attack steps show frontier AI’s research influence, with some demonstrating large-scale real-world effects.
Figure 2: A comprehensive categorization of frontier AI’s marginal risks to existing cyberattacks with qualitative impact assessment. Heatmap colors indicate impact levels: Green–no demonstrated effect, Yellow–demonstrated in research papers, Orange–demonstrated in real world, Red–large-scale real world impact. The figure reveals AI’s widespread application in achieving final attack goals but limited use in specific attack stages. This disparity exists because most reports disclose only final impacts rather than attack procedures, obscuring AI’s specific contributions. For precision, we include only reports with clear specifications for individual attack steps.
Attack step 1: Reconnaissance. Recent research explored frontier AI in multiple reconnaissance methods: active scanning [270], victim information gathering [73], and open-source database search [231]. In addition, real-world attacks increasingly use LLMs for active scanning (e.g., satellite capability understanding by Forest Blizzard) and information gathering about victims (e.g., by Salmon Typhoon) [200].
Attack step 2: Weaponization. Recent research demonstrated LLMs can generate functional Windows PE malware with high evasion rates against VirusTotal [213] ($>$70% for Ransomware, Worm, and DDoS). Real-world reports of AI-generated malware primarily focus on simple attacks like DDoS [23, 281]. Although LLMs are aligned to reject harmful queries, these alignment mechanisms can be evaded through jailbreaking [72, 187]. There are also early works on fine-tuning LLMs for attack generation, such as mutating malware to evade detection [248].
Research has leveraged LLMs for vulnerability identification across various network and software systems [253, 186, 174, 240, 80]. Real-world attacks have also begun employing LLMs for vulnerability detection, including MSDT vulnerability (CVE-2022-30190) [200] and SQLite overflow bugs [119]. Moreover, research showed LLMs can generate PoCs for memory corruption bugs in C/C++ [286, 311]. AI agent teams can exploit real-world N-day [90] and zero-day [91, 215] vulnerabilities (e.g., SQL/XSS injection, concurrency attacks, remote code execution) and solve CTF challenges [271, 329, 251]. However, the study on exploitation remains mainly at a research level, with only a few public real-world reports [142, 200, 295].
Attack step 3-5: Delivery & exploitation & installation. Initial access and installation have been demonstrated in the real world [208, 142] while persistence remains at a research level [245]. Initial access is achieved through compromised accounts or infrastructure [96, 208, 142], with AI assisting in password guessing [135, 288, 132] and bypassing defenses, including intrusion detection, CAPTCHA, and two-factor authentication [222, 74, 200, 70]. LLMs also assist malware configuration and installation [105]. Research works also used AI to maintain persistence across system restarts and credential changes, e.g., LLMs help masquerading by disguising malicious activities as benign [245, 28].
Attack step 6: Commend and control (C&C). Research demonstrates frontier AI can facilitate privilege escalation [121, 68, 158, 122]. For instance, ChainReactor [68] used LLM planning to generate exploit chains for privilege escalation. For stable and stealthy C&C channels, attackers can employ LLMs for automated domain generation, creating resilient remote management of compromised systems [245, 13, 28]. Recent research showed LLMs can assist with adaptation [57], allowing self-learning malware to create environment-tailored attacks. Despite these research advances, only a few real-world reports exist for credential access in this step [148, 244, 211].
Attack step 5: Action on objectives. As shown in Figure 2, AI-enhanced cyberattacks have proliferated across web, IoT, mobile, and cloud platforms, causing real-world financial losses, privacy leakage, data manipulation, and resource hijacking [23, 137, 281, 142, 96, 208, 63, 327, 99, 36, 316, 260]. A recent attack deployed AsyncRAT malware using LLMs to leak sensitive information [137]. Frontier AI has amplified attacks against web applications and IoT devices in eCommerce, including business logic abuse, API violations, bad bot attacks, and targeted DDoS on retail sites [191, 100]. These attacks have disrupted operations, stolen data, and exploited digital engagement for financial gain [142]. CrowdStrike reports a 110% rise in cloud-aware cyberattacks and 75% increase in intrusions year-over-year with AI enhancement [63]. For instance, threat actor INDRIK SPIDER exploited cloud-based credential storage, accessed Azure Key Vault credentials, and even used ChatGPT while managing Azure resources [63].
There are some early research exploring end-to-end attacks using LLMs or AI agents [306, 289, 280, 260]. SecCodePLT [316] used LLMs to generate end-to-end attacks with pre-specified attack steps and prompts. Autoattacker [306] proposed a multi-agent system with planning and generation agents to automatically analyze attack goals and generate corresponding attacks. They demonstrated that AI agents can plan and generate attacks for well-defined attack goals in a controlled environment.
Future impact. First, frontier AI could potentially lower the attack thresholds, particularly for attack steps 1&2, where it assists with information gathering even with limited access. Frontier AI models are trained on vast datasets, which potentially allows them to guess and reason partial information. For instance, deep learning models can achieve high accuracy in reverse-engineering stripped binaries [114, 220, 305]. In addition, as foundation models advance, they are expected to better understand program logic and semantics, likely enabling them to identify and exploit logical bugs under partial information (demonstrated by the emerging reasoning models [112, 203, 169]).
Second, we envision frontier AI can make attacks more automated and stealthy. For example, automated exploit chain generation in the attack step 3 (exploitation and installation) has been challenging. Previous approaches like counter-example-guided program synthesis [92] and rule-based methods [301] require expertise and significant manual effort for complex programs. Frontier AI offers solutions to overcome these limitations and other labor-intensive attacks, e.g., resource hijacking and operational disruption. Third, AI’s continuous learning capabilities enable automatic evolution and customized attack generation. By learning from both successful and failed attempts, attackers can refine strategies for more efficient and stealthy attacks.
[…]
                                         Attack step Coverage Metric Dynamic updates
[…]
3.1.2 Quantitative impact analysis
Table II summarizes existing benchmarks for evaluating LLMs and AI agents in different attack steps. In summary, except for vulnerability exploitation in step 2, existing benchmarks are limited in risk coverage, metric design, and dynamic evolution. Table VI in the Appendix details SOTA LLMs’ performances on existing attack benchmarks.
Attack step 1: Reconnaissance. Few benchmarks exist for reconnaissance. Most of them [274, 184, 286] evaluate LLMs through text-based reconnaissance questions and use another LLM as the judge. They fail to test the capabilities of generating and executing actual attacks. AutoPenBench [101] included reconnaissance but covers limited methods and scenarios. Results show that simple AI agents (LLMs with tool uses) demonstrate early success in reconnaissance tasks like discovering potential victims on remote networks or conducting surveillance. However, agents cannot identify target-specific vulnerable services due to limited service information from standard tools (e.g., Nmap) and the lack of deeper interactions with target systems.
Attack step 2: Weaponization. For malware creation, RedCode [111] benchmarked a prompt set that queries LLMs to generate malware in Python and uses VirusTotal as the judge. Another work crafted a benchmark to test the safety alignment of LLMs with jailbreaking prompts on attack generation [285]. They showed that Claude 3.5 Sonnet is more resilient than GPT-4o and Genimi-1.5. Regarding vulnerability exploitation, recent works designed some benchmarks for crash and PoC generation (e.g., CyberSecEval [286] and SecCodePLT [316]). They crafted specific coding tasks that may encounter known vulnerabilities in the CWE database and evaluate where LLMs generate vulnerable code when finishing these tasks. SecCodePLT [316] supports dynamic testing, which is more precise than LLM judgment. Their results showed that both SOTA commercial and open-source models frequently generate vulnerable code, with stronger models being more vulnerable due to their enhanced task understanding and execution.
While no dedicated exploit generation benchmark exists, existing CTF benchmarks [251, 329] and model pre-deployment tests [279] included sub-benchmarks evaluating exploitation capabilities. Their findings indicate SOTA commercial models like OpenAI-o1 and Claude 3.5 Sonnet achieve skill levels matching technical non-experts and cybersecurity apprentices in vulnerability exploitation. Claude 3.5 Sonnet can exploit more complex vulnerabilities at cybersecurity practitioner level. However, SOTA models rarely solve expert-level challenges, e.g., OpenAI-o1 has a very low success rate [329].
Attack step 3$\sim$7. Benchmarks for these later attack steps remain severely limited in coverage and quality. Similar to Reconnaissance, most existing benchmarks employ text-based queries rather than generating actual attacks [274, 184, 286, 156]. SecCodePLT [316] and AutoPenBench [101] include partial attack steps and use dynamic testing as the judge. SecCodePLT [316] provides a network system with pre-specified attack paths and targets, enabling end-to-end attacks, including reconnaissance, initial control, C2, and collection. Their results showed that Claude 3.5 Sonnet’s safety alignment rejects most attack generation queries, while GPT-4o generates end-to-end attacks with low success rates. CTF benchmarks [251, 329] include pwn challenges testing privilege escalation, showing that commercial LLMs slightly outperform open-source ones.
Future efforts. Existing benchmarks are limited across multiple dimensions. Constructing practical, comprehensive, high-quality benchmarks that can be dynamically updated requires significant community-wide effort. Section 7.1 provides actionable recommendations for achieving this goal.
Enhancing human attackers. Frontier AI not only enables fully automated attacks but also enhances human attackers’ productivity. At a high level, attackers can leverage AI to automate partial attack steps or provide initial attack solutions. For example, attackers can leverage AI to generate exploits for individual vulnerabilities and then chain them together manually. They can also further polish the LLm-generated phishing attacks to be more stealthy and convincing. While frontier AI will improve attack automation, we anticipate that in the short term, there will be more real-world attacks leveraging such human-AI collaborations.
Figure 3: Marginal risks of frontier AI to existing attacks against humans.
3.2 Attacks Targeting Humans
Humans remain the weakest link in security systems, with AI-driven attacks against them proliferating. Figure 3 shows that most human-targeted attacks have demonstrated large-scale impacts in the real world. Studies reveal a 135% increase in social engineering attacks following ChatGPT’s widespread adoption in 2023 [67]. Phishing attacks constitute a significant portion [43, 89, 30], with voice phishing surging 260% in 2023 compared to 2022 [262]. While studies from 2023 and mid-2024 [43, 88] found AI-generated phishing less effective than human-crafted attacks, recent research [128, 17] demonstrates fully AI-automated phishing can now match human expert effectiveness.
Malicious AI systems like FraudGPT and WormGPT are actively traded in underground markets [318], with FraudGPT selling for \$200 monthly or \$1,700 annually on dark web platforms and Telegram [243]. These systems analyze social media profiles and public data to craft tailored, convincing phishing messages [33]. AI-powered attacks have significantly increased identity theft and social manipulation [192, 206]. Companies now use AI to generate fraudulent advertising, fake reviews, and celebrity endorsements, prompting the US Federal Trade Commission to implement countermeasures against deceptive practices [258].
Deepfakes represent one of the most serious threats. Hong Kong police documented attackers using AI-driven deepfake technology to impersonate a chief financial officer during a video call, successfully defrauding \$25 million [49]. Child exploitation constitutes another severe AI misuse, including AI-generated child sexual abuse material, automated grooming, and detection evasion techniques [266]. AI-driven bots enable cyberbullying and harassment [134], facilitating emotional manipulation and behavioral control. Some nations reportedly employ AI for cyber espionage and psychological manipulation [238]. Misinformation presents another critical concern, with research demonstrating AI’s growing capabilities to influence human opinions and decisions [223, 21, 16, 104, 62]. These highlight how these capabilities can be exploited to spread misinformation and manipulate human behavior at scale.
Future impact. The impact of frontier AI attacks against humans will intensify dramatically. Advancing AI systems will enable more convincing, stealthy, and targeted social engineering attacks that bypass traditional security measures and human judgment. Multi-modal AI capabilities—combining text, image, audio, and video—will blur distinctions between authentic and fraudulent content, resulting in greater social damage from deepfake fraud and more effective public opinion manipulation through realistic synthetic media.
Key takeaways. Qualitative: Real-world AI-facilitated attacks have been reported across various systems, yet AI involvement in specific attack steps remains at an early stage and is limited to research settings. Steps 1&2 (reconnaissance and weaponization) show greater AI integration compared to later steps, with notable advances in malware creation and vulnerability exploitation.
Quantitative: Benchmarks show AI cannot generate valid malware but can exploit bugs at cybersecurity apprentice level. GPT-4o can generate end-to-end attacks, while Claude 3.5 Sonnet rejects most attack-related queries. Attack benchmarks can be improved in attack step coverage, metrics, and dynamic updates. Prediction: Frontier AI will lower the thresholds for attack steps 1 and 2 by reasoning with partial information, as well as enabling more automated and stealthy end-to-end attacks. Human: Frontier AI attacks already have a large-scale impact through social engineering, fraud, deepfakes, child abuse, and misinformation.
4 Frontier AI’s Impact on Defenses
This section discusses frontier AI’s impact on defenses for systems and humans, where we break down system defenses into empirical defenses and those with provable guarantees.
4.1 Proactive and Reactive Defenses
4.1.1 Qualitative analysis
Figure 4 shows our categorization of cyber defenses summarized from existing literature [261, 93, 124]. As shown in the figure, frontier AI has been predominantly utilized in proactive testing and attack detection but lacks widespread adoption in vulnerability triage and remediation. Currently, vulnerability remediation remains a labor-intensive and time-consuming process.
Defense step 1: Proactive testing include penetration testing and proactive vulnerability detections. Early research works automate penetration testing with LLMs and AI agents [155, 73, 120, 306, 273, 60, 133, 145] and more recent ones focused on optimizing agents’ workflows and tool calls [155, 73]. Due to the dual usage nature, these methods may also be used to facilitate attacks. LLMs have been used in industries for penetration testing [106] (But we lack statistics on whether they have been widely used).
More and more research leverages AI to facilitate vulnerability detection. This first includes developing customized graph neural networks [338, 307] and uses general LLMs and fine-tuned ones to directly identify vulnerabilities in various software applications [80, 5, 204, 240, 186, 174, 254, 224, 84, 269]. These works reported mixed results when comparing AI with traditional static analysis tools (e.g., CodeQL, Snyk, and Fortify) and are all concerned about the errors of AI and limited scalability to repository-level detection involving multiple functions and files. Early evidence showed that AI agents can help with vulnerability detection (e.g., Google Project Zero’s report on a real-world vulnerability: stack buffer underflow in SQLite) [327, 170]. Some research works also improved dynamic fuzzing with frontier AI [139, 253, 310, 207, 196, 76, 330], such as LLM-based pattern/rule/seed generation and RL-based power scheduling [255]. There are also early explorations in using LLMs for hybrid methods that combine static with dynamic program analysis, i.e., HyLLfuzz [195] developed an LLM-based concolic execution that uses LLMs as the solver for finding inputs to specific program branches.
Future impact. First, frontier AI could enhance program-analysis-based vulnerability detection approaches, better balancing soundness, completeness, and scalability. Specifically, AI agents’ scheduling and planning capabilities can improve dynamic fuzzing’s search efficiency and help select promising states for symbolic execution to prevent path explosion. These advancements can be extended to concolic execution and other hybrid approaches. However, frontier AI’s lack of robustness may offset marginal benefits, necessitating in-depth explorations that can design AI-based solutions for tackling traditional methods’ key challenges (see Section 7). Second, reducing AI-based vulnerability detection error rates requires improving models’ vulnerability reasoning capabilities. This requires high-quality datasets with additional reasoning data, such as vulnerable code with intermediate program states and human reasoning processes. Such supplementary information is crucial because frontier LLMs can process intermediate data to enhance decision-making beyond simple classification. Then, we need to design customized training approaches that improve the models’ reasoning capabilities, which help models handle ambiguous and complex cases (evidenced by recently emerged reasoning models). Third, with stronger models as agents, frontier AI could handle more complex cases for real-world code base.
Figure 4: Impact of frontier AI on different defense steps. It also shows the relations with the attack steps in Figure 2.
Defense step 2: Attack detection. Malware detection and network intrusion detection are two applications where ML has been widely used [290, 302, 45, 283, 161]. Recent research showed that frontier AI can address key challenges of traditional AI-based attack detections: reliance on feature engineering and labeled data, and poor generalizability. First, traditional AI models like decision trees and linear regression rely heavily on well-designed feature engineering, while deep learning can learn directly from raw inputs. When being used for attack detection, deep learning also eliminates the need for feature engineering and can explore large learning spaces beyond human-designed features, potentially leading to more effective solutions.
Second, existing AI-based attack detection requires large amounts of well-labeled data, which is challenging for security data [12, 302]. Transformers can learn general knowledge from vast unsupervised data [282, 77, 35], which can be customized for specific downstream tasks during supervised fine-tuning, significantly reducing reliance on large labeled datasets [82, 77]. Recent works showed that transformers trained on raw network traffic bytes outperform supervised models trained on handcrafted features in intrusion [117, 178] and malware detection [7, 324]. Third, existing AI-driven attack detection faces generalizability issues with out-of-distribution (OOD) data. Frontier AI models, trained on billions of data points or trillions of tokens, encompass wider distribution ranges during training. This reduces OOD likelihood during inference, which improves the generalizability in intrusion detection [117, 320, 98, 265] and malware detection [7, 324].
Future impact. Frontier AI-driven attack detections can continue improving in generalizability, particularly to attack evolvement and unseen threats, as well as adversarial robustness. As the models increase in capacity and reasoning capability, they could be better at finding hidden connections between existing attacks and evolved ones. In addition, through lifelong learning, these models can continuously update themselves to stay current with new threats. Furthermore, with enhanced safety alignment and more powerful and specialized guardrails, frontier AI models can become increasingly robust in threat detection. Recent advancements in model interpretability [340] allow developers to access a model’s internal representations and reasoning processes when a decision is made, which improves model interpretability and helps identify attacks. Finally, frontier AI could also be used for attack monitoring and instant alerting.
Defense step 3: Triage & forensic. First, for PoC generation, existing works explored improving dynamic fuzzing and symbolic execution with LLMs [292, 126]. For example, recent research [292] proposed using LLMs to translate Python path constraints into Z3 code, which enabled symbolic execution on Leetcode problems with complex control flows and list data structures. Second, for root cause analysis, traditional methods [114, 202, 64] employ static and dynamic analysis tools, such as reverse execution, backward taint analysis, value-set analysis, and alias analysis. Some recent works explored developing AI agents for root cause analysis [235, 308, 225]. These agents collect additional information (e.g., logs) through tool calling and leverage advanced prompting techniques (e.g., ReAct [317] and SelfReflection) to improve root analysis performance. Later in Section 4.1.3, we discuss how frontier AI facilitates SOC, which is also related to triage and forensics. Third, for reverse engineering, early works used recurrent networks for function boundary and signature recovery [257, 56]. However, such models have limited generalizability. Recent research [220, 305] explored pre-training or fine-tuning foundation models for binary codes and demonstrated better performance than RNN-based models and traditional rule-based tools in multiple reverse engineering tasks.
Future impact. Frontier AI can address key challenges in existing heuristic-based approaches for triage and forensic. For example, alias analysis is crucial for recovering data flow, especially with partial execution traces. Research demonstrates recurrent networks’ potential for alias analysis [114], and transformer-based models could further enhance this capability. Frontier AI can also enhance PoC generation and reverse execution as foundation models advance in capability, generalizability, and robustness. More broadly, we believe AI agents will become a promising approach that combines AI-based triage with traditional methods. Specifically, we can build multi-agent systems for root cause analysis, integrating existing techniques like reverse execution, taint analysis, and value-set analysis with AI-driven alias analysis as tools. In such a system, LLMs would drive the analysis by generating step-by-step plans, calling tools, and dynamically adjusting the analysis logic.
Defense step 4&5: Remediation development and deployment. Research demonstrates frontier AI’s capability to automatically generate security vulnerability patches [172, 177, 159, 3, 332, 197, 218, 263, 321, 55]. To enhance patching effectiveness, researchers have explored various prompt strategies and built AI agents where LLMs retrieve contextual information (e.g., caller functions of vulnerable code) [34, 177]. The AIxCC competition demonstrates real-world applications, including patching an Off-by-One bug in SQLite3 [119]. However, no specific work exists on AI-facilitated remediation deployment.
Future impact. We envision that specialized reasoning models can generate better patches for complex vulnerabilities. Early research showed that fine-tuning LLMs through RL or supervised learning can enhance bug-fixing capabilities for non-security bugs [309, 144]. We can also build more advanced agents integrated with traditional program analysis techniques to generate, validate, and refine patches. Recent research has demonstrated the effectiveness of this approach for logical and non-security bugs [171, 312, 304]. It is also critical to provide formal guarantees for LLM-generated patches, especially for security vulnerabilities [46].
Remediation deployment remains a lengthy process requiring substantial manual effort [42, 66], e.g., hospital systems currently average 471 days for remediation deployment [42]. As shown in Figure 4, new patches must undergo heavy testing to ensure reliability and correctness and need to resolve any potential conflicts with their dependencies. We envision that frontier AI can also accelerate and automate remediation deployment, given they can accelerate general software development (e.g., Copilot [103], Cursor [65]). For example, code generation models can generate testing cases for functionality and security testing [237, 54, 334]. LLMs can correct configuration errors (e.g., Hadoop) and set up software environments (e.g., Linux server) [250, 38]. As AI models continuously improve their ability to understand and analyze large code projects, they could be applied to identify potential conflicts in dependent packages and even fix such conflicts. These capabilities of frontier AI can largely accelerate the remediation deployment and reduce the time window during which systems remain vulnerable.
TABLE III: Overview of existing defense benchmarks. “Vul. D” and “Att. D” refer to vulnerability detection and attack detection. “RT” is root cause and “RE” is reverse engineering. “Re. dev.” and “Re. dep.” stands for remediation development and remediation deployment. The performances of the SOTA LLM models on these benchmarks are shown in Table VII.
                                      Defense steps Quantity Notable examples Key limitations
1 Pen. test Few AutoPenBench [101] Limited attacks, target systems, metrics,
Vul. D Many PrimeVul [80],FuzzBench [199], OSS-Fuzz [247] limited complexity and diversity, low-quality labels
2 Att. D Many IDS2018 [252], Drebin [19], BODMAS [315] Low-quality data and labels, lack OOD/hidden tests
3 PoC & RT Few CRUXEval [108] Low complexity, no benchmark for root cause,
Rev. Eng. Many ByteWeight [25], ReSym [305] data leakage, incomplete task coverage
4 Rem. Dev. Few SWE-bench [147] Noisy, no benchmark for security bugs, limited programming languages (PLs)
5 Rem. Dep. None None N/A
4.1.2 Quantitative analysis
Table III shows an overview of existing defense benchmarks, where most benchmarks focus on vulnerability detection and attack detection, while other defense steps lack high-quality benchmarks. Table VII details SOTA LLMs’ performance on these benchmarks.
Defense Step 1: Proactive testing. Benchmarking penetration testing is related to attack benchmarking. As discussed in Section 3.1, penetration testing benchmarks remain at an early stage, where AutoPenBench [101] is the most recent benchmark enabling end-to-end penetration testing with dynamic metrics.
There are many vulnerability detection benchmarks, including static benchmarks with labeled code snippets [217, 80, 277, 47, 275, 284, 118, 212] and dynamic benchmarks with fuzzing instrumentation [247, 199]. The latest static benchmark, PrimeVul [80], contains vulnerable C/C++ and Python code covering over 100 CWEs, showing that SOTA LLMs achieve only $\sim$30F1 score. PrimeVul shares limitations with other static benchmarks, most notably low-quality labels. These include inaccurate labels caused by missing context (snippets only vulnerable under specific conditions) and incomplete labels (snippets with multiple vulnerabilities but only one labeled). In addition, existing benchmarks mostly cover snippets with only one function selected from limited projects. For dynamic datasets, OSS-Fuzz [247] is the largest fuzzing benchmark, and it is actively maintained and updated. FuzzBench [199] is a smaller but less noisy benchmark widely used in research papers to test novel fuzzing methods. These benchmarks are also used to test AI-facilitated fuzzing approaches, which achieved improvements over traditional fuzzers [256].
Future efforts. For penetration testing, future works should expand attack step coverage, diversify target systems, and implement more comprehensive evaluation metrics. For vulnerability detection, we should enhance label quality through human validation and include complex code snippets from diverse projects.
Defense Step 2: Attack detection. Many benchmarks exist for network intrusion detection and malware detection [252, 19, 315]. Recent research [18] highlights these benchmarks’ limitations in data quality (duplicated data, shortcuts) and label accuracy. Foundation model-driven detection shows very high accuracy on these benchmarks [117, 178]. Besides attack detections, CTIBench [8] designed tasks for LLMs to pinpoint the specific CWEs and attack techniques for given attack instances
Future efforts. Future work can improve data and label quality while incorporating more complex attack instances. Since existing AI-based detection lacks generalizability, benchmarks can include OOD testing sets with unseen attack types and drift patterns. Moreover, dynamic benchmark updates are essential to incorporate emerging attacks.
Defense Step 3: Triage & forensic. CRUXEval [108] benchmarks PoC generation for vulnerable C/C++ codes with overflow bugs, showing that GPT-4o achieves 75% pass@1 success rate.
Future efforts. We can include more vulnerabilities and cover more programming languages for PoC generation benchmarks. While no benchmark is publicly available for comprehensively evaluating AI-driven reverse engineering, most AI-based methods released their own datasets [257, 114, 305]. These datasets could potentially be merged to create a unified benchmark, covering tasks from function boundary identification to variable or data structure recovery. There is no public benchmark for root causes, and constructing such benchmarks is challenging because ground truth is hard to determine. One possible way is to extract samples from reproducible vulnerability datasets (e.g., ARVO [193]) and craft ground truth root causes via crowdsourcing.
Defense Step 4&5: Remediation development and deployment. Given that AI is not yet widely involved in these steps, there are not many benchmarks. The most popular benchmark is SWE-bench [147], which extracts issues from multiple Github Python projects. To fix the issues, we need to build AI agents that can analyze and locate the issues, generate patched code, and validate the generated patches. SOTA multi-agent systems with OpenAI o3 and Claude 3.5 Sonnet can resolve more than 55% issues in the SWE-bench-verified benchmark. Recent works extended the SWE-bench to multi-modal settings [312] and the Java language [325].
Future efforts. Given that the SWE-bench focuses on non-security bugs, it is necessary to construct benchmarks with security bugs covering multiple languages, especially C/C++. The procedure can be similar to root cause where we start from datasets with reproducible vulnerabilities. The main effort will be creating execution environments, crafting ground truth patches, and defining proper evaluation methods (e.g., writing test cases for dynamic testing).
4.1.3 Facilitate human defenders and SOC
Frontier AI can also improve the productivity of human defenders. Human and AI collaborative defenses could draw a balance between defense efficacy, robustness, and cost efficiency. For example, during proactive testing and attack detections, we can use AI to filter out the most obvious benign cases, which significantly reduces human efforts. AI can also propose initial PoCs, patches, and configurations, such that a human only needs to examine and fix these candidates rather than starting from scratch. Frontier AI is also enhancing Security Operations Centers (SOCs) by improving threat detection, incident response, and operational efficiency [95, 83]. Recent progresses [116, 259] show that AI-driven tools can autonomously investigate and respond to security alerts, enabling SOCs to analyze increasing amounts of security data and identify more threats in real-time. The continuous advancements in AI-driven tools are expected to enhance the productivity and scalability of SOCs [157].
4.2 Defenses with Provable Guarantees
Secure-by-design and secure-by-construction defenses provide formal security guarantees. Formal verification proves whether systems satisfy specific properties, with tools developed for various systems [246, 109, 164, 125, 48, 31], including seL4 [246] and CertiKOS [109] for OS kernels, CompCert [164] for compilers, IronFleet [125] for distributed systems, and EasyCrypt for cryptographic proofs [27]. This process is labor-intensive; seL4 required tens of proof-engineer years [129, 130]. Failed verifications require expert refinement, and software updates demand rewriting proofs, significantly increasing human effort.
Recent research explores frontier AI to improve formal verification’s automation and scalability. Some works enhanced interactive theorem proving [138, 322, 10, 314], e.g., GamePad [138] uses deep learning to synthesize Coq proofs. Others focus on automated program verification, where the typical solution is to model programs in domain-specific formal languages (DSLs), formulate security properties as specifications, and use constraint solvers for verification. Research works used LLMs to generate program invariants, lifting a program to a DSL or supplementary text hints [219, 150, 32, 300, 185]. For example, Loopy [150] leveraged LLMs to generate candidate loop invariants and refined invariants via an SMT solver. Lemur [300] used LLMs to generate high-level program invariants as sub-goals, which are then checked by solvers. LLMLift [32] combined LLMs with verified lifting tools to provide verified code transpilation from the source to DSLs, where LLMs generate lifted codes and proofs for functional equivalence. DafnyBench [185] evaluates SOTA LLMs in generating hints for the Dafny verifier. Other works improve solvers, including using traditional deep learning to enhance SAT and SMT solvers [24, 173] and leveraging LLMs for solver testing and fuzzing [268]. Section 7.2 shows our recommendations for improving formal verification with frontier AI.
4.3 Defense For Humans
Frontier AI enhances defenses against human-targeted attacks. LLMs show promise in detecting social bots [94], and AI has been deployed for real-world fraud detection [166, 326, 78]. Research demonstrates multi-modal agents capable of detecting phishing websites [39]. AI-driven solutions are increasingly being applied to real-world scenarios. For example, one company deployed an AI-based honeypot to combat vishing, using a chatbot with a “grandma” voice to occupy fraudsters and waste their time [198]. AI-based deepfake detection continues as an active research area [229], with exploration of LLMs’ capabilities in identifying fake images [6]. Another significant direction focuses on AI-powered misinformation detection [14, 79, 180], often employing LLMs to identify LLM-generated content. Watermarking techniques represent another promising detection approach [333, 214, 179, 335, 110].
Despite these active research efforts, Figure 3 shows current defensive techniques struggling to match sophisticated attacks. Research reveals significant limitations in AI systems’ defensive capabilities. For example, studies examining GPT-4’s ability to detect fake images [6] concluded the model lacks reliability and shows limitations with complex tasks. While many papers propose AI-powered defenses against deepfakes and social bots, most address traditional attack scenarios rather than novel or sophisticated attacks enabled by frontier AI. We need studies on adversarial dynamics between defensive and attack AI systems—for example, how defensive AI can detect deepfakes or social bots generated by attack AI. This gap underscores the urgent need for research developing AI-driven defenses against AI-driven human-targeted attacks.
Future impact. As frontier AI evolves, its defensive capabilities will improve significantly. Future systems can detect and prevent attacks in real-time by analyzing patterns across multiple modalities simultaneously. These advanced systems could predict and preempt sophisticated social engineering attempts, deepfakes, and fraud schemes before reaching their targets. In addition, an increasing number of AI systems will be deployed as decoys [198, 230], posing as vulnerable targets to deceive attackers and exhaust their resources. By integrating frontier AI, authentication mechanisms can become more robust and adaptive to emerging threats, including sophisticated watermarking techniques and detection methods designed to counter evolving attack strategies.
However, this defensive potential faces a critical challenge: the inherent arms race between offensive and defensive capabilities suggests that as defensive systems become more sophisticated, attackers will likely develop increasingly advanced countermeasures using the same frontier AI technologies. We discuss this arms race between attacks and defenses in the next section.
Key takeaways. Qualitative: Frontier AI has been utilized in proactive testing and attack detection but remains underexplored in vulnerability triage and remediation. Existing AI-driven detections suffer from high false positive rates, limited generalizability, and insufficient robustness. Early AI-driven triage and patch tools cannot beat heuristic-based methods and manual analysis, nor can they provide formal correctness and security guarantees. Quantitative: Benchmarks show frontier AI outperforms traditional static detection methods and modestly improves fuzzing for vulnerability and attack detection. Future benchmarks for proactive testing and attack detection can enhance data quality and label accuracy, and include hidden OOD test sets. More comprehensive benchmarks are needed for other steps. Prediction: By training with high-quality datasets with rich program artifacts and stronger safety alignments, AI-driven detections can have better reasoning capabilities and improvements in generalizability and robustness. AI agents, integrating AI-driven planning and generation with program analysis tools, will harness the strengths of learning-based and program analysis-based approaches, potentially leading to more effective and automated triage and remediation. Systems with provable guarantees: Early attempts leveraged frontier AI for theorem proving and program verification. Frontier AI advances formal verification by automating theorem proving, generating invariants and specifications for complex systems, and improving solver efficiency and scalability. Humans: Frontier AI demonstrates limited capability in mitigating sophisticated human-targeted attacks (e.g., deepfakes), necessitating intensified research in this area.
5 Will Frontier AI Benefit Attackers or Defenders More?
Building on our analysis of AI’s impact on cyber attacks and defense, this section predicts whether frontier AI will likely benefit attackers or defenders more in the short term. We base our analysis on: 1) equivalence class, where new defenses may equally benefit attackers, 2) the fundamental asymmetry between attackers and defenders, and 3) economic perspectives.
TABLE IV: Equivalence classes: A list of defense and normal capabilities that can also help attacks.
  Defense stage Defense capabilities Attack usages
Proactive testing Pen. testing Enable more targeted attacks
Vulnerability detection Find vulnerabilities in target systems
Attack detection ML-based threat detection Develop stronger evasion methods
Lifelong monitoring Re-purpose it to monitor defenses
Triage forensic PoC & root cause Facilitate localization & exploitation
Reverse engineering Understand targets and steal source core
Remediation Patch & testing generation Malware & weapon & exploit generation
Automated configuration Automated installation and gain access
Normal utilities Multimodal generation Automated reconnaissance and delivery
5.1 Equivalence-Class-Based Prediction
In cybersecurity, many defensive tools function as double-edged swords usable by attackers. This applies equally to frontier AI-enabled capabilities; AI advancements in software engineering and defense can help attackers develop malware, command-and-control systems, and botnets. These are formally termed equivalence classes, where defensive capability development inherently benefits attackers.
Table IV shows equivalence classes across the defense lifecycle in Figure 4. Techniques developed for penetration testing, if obtained by attackers, enable more targeted attacks during reconnaissance and weaponization. Since penetration testing methods are system-specific, they provide attackers with customized attack tools. Attackers can also repurpose vulnerability detection tools to identify system weaknesses. As AI models develop stronger memorization and reasoning capabilities, they can potentially identify vulnerabilities with limited information, lowering attack thresholds for attackers with restricted target access.
Triage and forensic techniques create additional equivalence classes. AI-driven PoC generation and root cause analysis benefit attackers for vulnerability localization and exploit development. Reverse engineering techniques give attackers deeper target insights, potentially including source code. As AI-driven reverse engineering improves, particularly with stripped binaries and obfuscation, attackers gain additional advantages.
With frontier AI’s improved code understanding and generation, patch and test-case generation capabilities grow—but these same capabilities help attackers generate weapons and chain exploits, creating large-scale threats. If agentic systems automate patch deployment, they can be repurposed to deliver and install malware. Once remote control is established, generation and deployment-capable agents can potentially manage later attack stages, including target location, action execution, and exfiltration.
Beyond defense capabilities, frontier AI’s general utilities serve malicious purposes. As discussed in Section 3.2, multimodal capabilities of frontier models enable the creation of sophisticated deepfakes.
5.2 Asymmetry-Based Prediction
Beyond equivalence classes, inherent attacker-defender asymmetries give attackers disproportionate advantages when using frontier AI, stemming from three key factors.
Asymmetry in cost of failures. A clear asymmetry exists between attackers and defenders in their goals and requirements. Attackers need just one successful exploit, while defenders must guard against all possible threats and vulnerabilities [261]. This fundamental difference severely constrains defenders’ failure tolerance. For AI-driven threat detection systems, both false positives and false negatives carry serious consequences: false positives disrupt normal functionalities, while false negatives leave threats uncovered. Even a single false negative can compromise the entire defense. Similarly, for automatic patching, ensuring accuracy before deploying machine-generated code is essential, raising capability requirements and necessitating additional verification processes that inevitably require human intervention. Compounding this issue, frontier AI’s inherent probabilistic nature introduces errors that further restrict its deployment in critical defense systems.
Conversely, attackers have much higher error tolerance when adopting new technologies. Failed attacks can be rerun or adjusted with minimal consequences. This enables attackers to exploit emerging technologies more readily, even those still in early development stages.
Asymmetry in remediation deployment and required resources. As discussed in Section 3.1, deploying remediation is time-intensive even when effective. The process requires testing for correctness, resolving dependency conflicts, global deployment, and post-deployment verification. This labor-intensive procedure demands significant resources. Legacy systems take even longer to patch, with many remaining perpetually vulnerable. Conversely, attackers can operate with minimal resource requirements. They can monitor vulnerability patches, generate exploits using public information about legacy systems, and target unpatched systems. They exploit both permanently unpatched systems and the deployment time gap between patch development and complete rollout.
Different priorities of scalability and reliability. Attackers prioritize scalability, while defenders emphasize reliability. This makes AI a natural choice for attackers, as its primary capability is reducing human involvement and increasing automation. AI lowers barriers for attackers by decreasing requirements for deep expertise, specific target system knowledge, and execution time and effort. Even if an attack fails against one system, it may succeed against another. For example, in large-scale attacks, attackers may simultaneously target thousands or millions of systems (e.g., Mirai Botnet [58]). Conversely, frontier AI models face reliability limitations in robustness, transparency, and generalizability. When used for malware detection, users cannot understand why specific software is flagged as malicious or predict when false predictions will occur. This reliability deficit hinders defenders from fully trusting and deploying AI techniques in critical defense systems.
This asymmetry may be magnified in critical physical infrastructure, where attackers exploit these infrastructures and propagate the damage to other systems. Beyond the asymmetry that favors attackers, certain forces and security paradigms can help shift this imbalance. For instance, large organizations typically command far greater computing resources and security expertise than individual attackers. By leveraging centralized, scalable infrastructure, they can implement comprehensive, multi-layered defenses as part of a defense-in-depth strategy. Additionally, secure-by-construction paradigms and robust defense mechanisms—such as memory safety, control flow integrity enforcement, and privilege separation—offer strong guarantees against specific vulnerability types, substantially limiting attackers’ capabilities. While these approaches help reduce attackers’ advantages, fully realizing them will likely take time.
5.3 Attack-Defense Dynamics From the Economic Perspective
Economic rationality drives both attackers and defenders—attackers strike only when expected profits exceed costs, while defenders implement protections only when benefits outweigh costs. As we have seen in Sections 5.1 and 5.2, frontier AI is changing these economic considerations significantly. By enabling more frequent and stealthy attacks, AI increases defensive costs while potentially rendering traditional defenses obsolete. This cost imbalance may force resource-constrained defenders to abandon certain protective measures and attackers will achieve higher investment returns due to diminished defenses. Appendix .2 shows a formal mathematical analysis of economic dynamics.
5.4 Evolutionary Attack-Defense Dynamics
The diverse resource and capability spectrum across attackers and defenders shapes how frontier AI impacts attack-defense dynamics over time. Attackers range from individual actors with limited resources to sophisticated state-sponsored groups with extensive funding. Similarly, defenders span from small organizations with basic security measures to large enterprises deploying advanced defense systems. Frontier AI lowers barriers to successful attacks by reducing resource and expertise requirements, likely increasing attack frequency and scale. In the early phase, well-resourced attackers amplify their capabilities by developing specialized AI models and agents, enabling more sophisticated, stealthy attacks with large-scale societal impact.
Resource-constrained defenders see limited short-term benefits from AI adoption. As detailed in Section 4.1, remediation and deployment remain time-intensive, laborious processes. Even with AI enhancing attack detection, overall defense effectiveness is constrained by these operational limitations; resource-constrained defenders struggle to leverage frontier AI to address these operational issues due to limited resources and expertise. Well-resourced defenders can leverage frontier AI to address fundamental challenges, but this requires more time. These factors indicate attackers gain advantages in the early evolutionary cycle.
Moreover, AI’s inherent randomness and the bug-finding process introduce another factor affecting dynamics. One hypothesis suggests that due to AI outputs’ statistical nature and exploration randomness, defenders and attackers may discover different vulnerabilities when analyzing the same systems. This discovery divergence creates a critical advantage for attackers, who can exploit vulnerabilities they identify while defenders address an entirely different set, even when investing more resources in bug finding. This misalignment could significantly favor attackers.
Frontier AI capability advancement speed adds another dimension. Rapid AI capability acceleration could enable numerous vulnerability discoveries within short timeframes. This further tilts the evolutionary trajectory toward attackers, as defense deployment takes longer than attack execution. While attackers quickly leverage newly discovered vulnerabilities, defenders remain constrained by inherent delays in implementing patches across infrastructure. This timing asymmetry creates a pattern where attack capabilities surge ahead during rapid AI advancement periods, particularly benefiting resource-rich attackers who quickly and effectively adopt cutting-edge techniques.
Long-term, however, the evolutionary trajectory may gradually shift toward defenders as several counterfactors emerge. First, defenders could gain priority access to advanced techniques (e.g., OpenAI’s GPT-4o early access for safety researchers). As frontier AI advances, defenders benefit from improved automation, transitioning from manual patch development to AI-generated patches with automated verification. Second, as defenders use AI to systematically discover and patch vulnerabilities, systems become increasingly robust. In particular, by leveraging AI for formal verification and adopting a secure-by-design approach, defenders can build fundamentally more secure systems with provable guarantees. This continuous security improvement creates an environment where finding new vulnerabilities becomes significantly more challenging and costly for attackers. Once defenders achieve stronger system security, attackers must invest substantially more resources to find attack vectors, potentially making many attacks impractical or cost-prohibitive. Lastly, as regulatory frameworks and law enforcement mechanisms mature, attackers can face more severe penalties once caught, substantially raising failure costs and potentially deterring malicious activities.
Frontier AI will benefit attackers more than defenders in the short term. Reason 1: Frontier AI functions as a double-edged sword in cybersecurity — defensive capabilities can be repurposed for attacks (see Table IV). Reason 2: Attacker-defender asymmetry favors early AI adoption by attackers. Attackers need only succeed once, while defenders must guard against multiple threats with lower failure tolerance.
Economic perspective: Frontier AI shifts cybersecurity economics by reducing attack costs while increasing effective defense expenses, potentially forcing resource-limited defenders to abandon certain protections, though this dynamic may eventually reverse in defenders’ favor in the long term. Evolutionary dynamics: Multiple factors influence attacker-defender dynamics, including resource/capability spectrums, AI output variability, and AI development speed. Initially favoring attackers, this balance may gradually shift toward defenders over time.
6 Hybrid Systems and Security Challenges
AI agents & hybrid systems. As mentioned in Section 1, AI agents that combine LLMs with non-ML system components can enhance LLM-generated content quality and automate complex tasks. A simple example is Retrieval-Augmented Generation (RAG), which is also adopted in security applications [227]. They allow LLMs to reference external information and reduce hallucinations [167]. Beyond RAG, recent works built web agents [75, 336] for human-web interactions, coding agents [65, 312] for code writing and debugging, and personal assistants [242, 210] for daily tasks like calendar management and email. Recently, OpenAI developed deep research, an agent for conducting research tasks [209]. The security community has begun constructing multi-agent systems for complex attacks and defenses [235, 304]. An example design is to use a planning agent to decompose tasks into sub-tasks, with specialized agents collaborating with non-ML components to complete each sub-task. These examples show a shift toward increased AI integration into traditional software.
Marginal risks of hybrid systems. In Section 3.1, we identify two types of frontier AI marginal risks: augmenting existing cyber attacks and introducing new risks to hybrid systems. Figure 5 shows our taxonomy of new risks in hybrid systems, which fall into two categories: risks in foundation models and risks in AI agents/hybrid systems.
First, foundation models are vulnerable to both training-phase and testing-phase attacks. Based on existing literature [287, 267, 205, 41, 40], we outline several predominant risks that have been demonstrated in real-world commercial models. We also summarize existing automated red-teaming methods, particularly for testing-phase risks. In real-world practices, the predominant approach is to combine automated red-teaming with human red-teaming [272, 2, 276]. Attack prompts are fed to the models either directly or through (in)direct prompt injection, where attack prompts are appended after benign user prompts [183, 181].
Second, for hybrid systems, risks become more complex due to interactions between AI and symbolic components, which can be broadly classified into two categories: attack the AI component through other components and attack traditional symbolic components through AI components. Indirect prompt injection is a typical example of the first type, where attack prompts are embedded in other components rather than directly input into the model [69, 107, 341]. An example of the second type is an attack against an agent with a code generation model, a Python interpreter, a web server, and a database, designed to automatically handle data queries with proper access control. Here, an attacker might prompt the model to generate a privilege escalation code, granting illegal access to sensitive information [299]. Similarly, when an AI-driven chatbot is used in a customer service system, attackers can potentially trigger the chatbot to output inappropriate content that leads to a wrong refund [37]. Due to their complexity, solely crafting attack prompts is insufficient for effective red-teaming for hybrid systems. Recent work has started designing AI agents for red-teaming hybrid systems, mainly web agents [297, 296, 175], but this is still in its early stages. In the future, we believe there will be more automated red-teaming that works together with human-based red-teaming.
Figure 5: Taxonomy and red-teaming approaches of AI-augmented hybrid systems’ marginal risks.
Defenses for hybrid systems. Most of the existing defenses target models themselves with a focus on defending against jailbreaking attacks and prompt injection attacks. Technically speaking, such methods either conduct adversarial retraining of the target model [319, 51] or construct additional guardrail components for input or output filtering [323, 143, 182, 152]. There are only a few early explorations of defenses for hybrid systems [303, 298, 216, 339]. These methods primarily defend against attacks during tool calling, either by proactively adding data or control flow constraints[303, 298], detecting and blocking malicious tool calls [339], or implementing post-facto validation with undo features and damage confinement[216]. As hybrid systems become increasingly popular and complex, more advanced and comprehensive defenses will be in great need.
There is no work on designing defenses against hybrid systems with provable guarantees. It is a challenging task for the following reasons. First, existing formal verifications are designed for traditional symbolic programs. This makes them difficult to apply to non-symbolic programs, such as deep neural networks, which lack precisely specified properties and goals. For example, it is difficult to formally define the pre- and post-conditions for a real-world self-driving controlling system. While some early explorations exist, they lack stability and can only verify small models [154, 291]. Second, numerous studies have developed certified defenses for deep learning models [59, 165], and recent work has extended these approaches for certified defenses against jailbreaking attacks on LLMs [160, 233]. However, these efforts are still in their early stages and do not yet provide certified defenses against AI agents. Given their significance, further research is crucial to developing provable guarantees for hybrid systems.
Key takeaways. Hybrid systems: Frontier AI will drive the emergence of hybrid systems that integrate symbolic components and non-symbolic AI components. Risks: Frontier AI will introduce new marginal risks to hybrid systems at the model and system level (specified in Figure 5). Defense: Existing defenses for hybrid systems mainly focus on model-level, comprehensive system-level defenses are still lacking but will become increasingly necessary and important. None of the existing provable defenses, whether for traditional symbolic systems or machine learning, can be directly applied to hybrid systems.
TABLE V: Summary of our call to action to the security community and AI developers for improving and leveraging frontier AI to strengthen cybersecurity.
     Priorities Directions       Summary
  Marginal risk assessment Benchmark risks in existing attacks
Current
status
  Lack benchmarks for fine-grained attack steps that evaluate whether AI can generate executable attacks beyond texts
  Lack evaluation platforms with accurate metrics rather than LLM judgment
  Lack of attack benchmarks with carefully designed prompts for LLMs and comprehensive tool sets
Call-to-
actions
  Build benchmarks for fine-grained attack steps that cover rich target systems and possible attack paths and methods
  Combine benchmarks for individual steps to test whether AI can generate end-to-end attacks
  Involve necessary human efforts to ensure the environment correctness
  Ensure the data quality, e.g., design prompts that can possibly bypass the safety alignments of LLMs
  Design comprehensive evaluation metrics that cover both static and dynamic metrics
  Include non-ML tools that are utilized by attackers to support evaluating agent-based attacks
  Dynamically updates the benchmarks to include new attacks and support new AI tools
Benchmark new risks in hybrid systems
Current
status
  Lack benchmarks and advanced red-teaming methods for emerging models
  Lack risk categorizations and benchmarks for hybrid systems
  Lack automated red-teaming methods for hybrid systems
Call-to-
actions
  Evaluate existing model-based risk benchmarks on reasoning and small models
  Design specific benchmarks for reasoning and small models if needed
  Categorize risks for different hybrid systems and build benchmarks accordingly
  Design agent-based red-teaming methods for hybrid systems under realistic threat models
  Enhance cyber defenses Improve proactive testing
Current
status
  AI-driven penetration testings are far from comprehensive
  Program analysis-based detections still suffer key limitations in soundness, precisions, or scalability
  AI-based detections have limited generalizability to new vulnerabilities and scalability beyond the function level
  Existing benchmarks lack comprehensiveness in penetration testing and have noisy data
Call-to-
actions
  Build penetration testing benchmarks with more comprehensive targets and tool supports
  Design penetration testing agents with stronger planning under comprehensive tool sets
  Enrich static analysis-based detections with rules extracted from AI models
  Improve static analysis-based detections in state pruning with LLMs and RL
  Improve fuzzing in seed&grammar generations with LLMs and seed&mutator scheduling with RL
  Fine-tune reasoning models that better reason program states, dependencies, and vulnerabilities
  Fine-tune models with long context limits or build AI agents to scale AI-driven attacks beyond functions
  Build vulnerability detection benchmarks with correct labels, annotations, balanced distributions, and multiple levels
Enhance AI-based attack detection
Current
status
  Detect unseen attacks are still challenging
  AI-based detections have limited robustness and explainability
  Benchmarks have data and label problems, such as spurious correlations, biased samples, wrong and missing labels
Call-to-
actions
  Construct high-quality benchmarks that avoid the common data and label issues and dynamically update the datasets
  Design customized learning methods to address security-specific challenges, such as attack shifts
  Improve model robustness with safety alignments and release the model with interpretability tools
  Combine AI-based detection with traditional methods to balance error tolerance, robustness, and scalability
Automate triage & remediation
Current
status
  Lack automation in triage and remediation deployment
  Automated patching lacks scalability and correctness
  Lack high-quality benchmarks for these defense steps
Call-to-
actions
  Build multi-agent systems for root cause analysis with complex planning and program analysis tools
  Build patching agents with more accurate and comprehensive validation through differential fuzzing
  Build patching agents with formal security and correctness guarantees through symbolic executions and model checking
  Enable the agents to digest feedback from program analysis tools and iteratively refine its outcomes
  Fine-tune code reasoning models for PoC, patch, and secure code generations based on agents’ feedback
  Fine-tune specialized small models to replace the planning LLMs in triage and patching agent
  Leverage AI for automated deployment, e.g., testing case generation and fixing conflicts
  Construct benchmarks for more scenarios and program languages in triage and patching
Automate provable defenses
Current
status
  Formal verifications (FV) are labor intensive and lack scalability
Call-to-
actions
  Train and use AI to generate complex proofs of theorem proving
  Train and use AI to generate inputs for program verification (e.g., invariants and function summaries)
  Improve solver efficiency with AI-based planning
  Build secure hybrid systems Design novel security mechanisms
Current
status
  Lack real-time attack detection and monitoring for hybrid systems
  Lack proper system protection mechanisms, traditional protections cannot be directly used to hybrid systems
  Lack defenses with provable guarantees for hybrid systems
Call-to-
actions
  Design efficient monitors for both AI and symbolic components and periodically update them
  Develop system- and component-level security protections for hybrid systems (e.g., generalize privilege isolation)
  Build efficient verification for AI components in the hybrid systems and the entire system as a whole
  Model developer & users Model testing & transparency
Current
status
  Existing pre-deployment security testing mainly focuses on solving CTF challenges
  Lack of transparency and robustness
Call-to-
actions
  Conduct more comprehensive pre-deployment cyber testing for broader AI products
  Provide explanations to AI decisions and disclose certain training information
  Design comprehensive blue-teaming under worst-case scenarios, such as adversarial training and additional guardrails
Protect humans & user
Current
status
  The AI-powered attacks have impacted humans on a large scale
  The development of defenses lags far behind attacks
Call-to-
actions
  Develop AI-powered defenses against malicious social bots
  Implement AI-driven educational systems to enhance user awareness
7 Call to Action
In this section, we provide concrete recommendations for steering frontier AI toward strengthening cybersecurity. We focus on five priorities: risk assessment as the critical first step in developing effective countermeasures, leveraging frontier AI to strengthen defenses, designing secure hybrid systems, enhancing AI model development practice and transparency, and mitigating human-related risks. Table V summarizes our recommendations.
7.1 Continuous Marginal Risk Assessment
Our first call to action is for the security community to collaborate in constructing risk assessment benchmarks. This requires covering the fine-grained risk categories in Figures 2 and 5 to enable a comprehensive assessment of AI’s marginal risks. It also demands continuous tracking of real-world AI-powered attacks and benchmark updates to reflect evolving risks. Below, we provide concrete recommendations for constructing benchmarks to assess risks of existing cyberattacks and emerging hybrid system risks.
Build systematic benchmarks incorporating fine-grained attack steps. As discussed in Section 3.1, existing attack benchmarks have deficiencies in risk coverage, data quality, and metrics. To improve risk coverage, one can leverage existing cyber ranges like MITRE’s Caldera [61] and IBM Cyber Range [141], which provide environments for various attack steps, particularly steps 3$\sim$7 where benchmarks are lacking. We can adapt these human-oriented environments with AI interfaces, tool calls, and evaluation metrics, while dynamically adding more fine-grained attack steps. Constructing attack benchmarks across diverse target systems, including private networks with different endpoints, is also essential.
For challenging tasks like CTFs, we need to incorporate tool calling capabilities and support agent construction frameworks (e.g., LangGraph [162]) to evaluate different agent designs and identify effective tools. Finally, we need to integrate existing non-AI attacks and defenses to quantify AI-specific risks, measuring both fully automated attacks and human uplift—how AI enhances human attack capabilities.
Build benchmarks and red-teaming approaches for new risks in hybrid systems. As discussed in Section 6, existing benchmarks on AI safety primarily focus on model-level risks. With the emergence of small and reasoning models, testing and extending current model-level risk benchmarks to these new models is valuable. More importantly, the community lacks risk classification and fine-grained benchmarks for hybrid systems. We must conduct fine-grained risk classification for hybrid systems and build specialized benchmarks for each category.
Since hybrid systems are highly specialized, different types will present unique risks. We recommend first categorizing hybrid systems (e.g., web agents, coding agents, personal assistants as described in Section 6) and identifying their associated risks under realistic threat models. As discussed in Section 6, we can classify risks based on their targets: AI components or symbolic components. Currently, the predominant risk is through indirect prompt injection, attacking AI through symbolic components (tools). However, there were some early attacks targeting symbolic components through AI, such as AI-generated exploits against coding agents. This underscores the critical need for thorough risk evaluations of hybrid models.
Existing red-teaming for hybrid systems only explored limited risks of specific agents (e.g., interpreter abuse of code agents [286] and prompt injection of web agents [175]). Developing automated red-teaming frameworks for broader hybrid system risks is necessary, such as assessing multi-agent system risks through game theory or designing new attacks for systems with multi-modal agents.
Continuously ensure quality and realism of benchmarks. Creating effective cybersecurity benchmarks requires careful design since agents are tasked with exploiting vulnerabilities. When benchmarks contain flaws, agents can "cheat" by completing the cyberattack task through unintended pathways. Given the complexity of cybersecurity tasks involving various files and servers, ensuring benchmarks are properly built and solvable is crucial. For instance, Transluce noted that several of the Intercode-CTF benchmark tasks are broken or unsolvable [194, 313].
Moreover, human expertise remains essential for creating quality environments and data. Proper execution environments ensure AI explores legitimate attacks rather than exploiting environment flaws (as when OpenAI-o1 hacked a CTF’s Docker container [146]). Research shows that the automated collection of vulnerable code commits without human oversight often yields non-vulnerable or security-unrelated samples [277]. Balancing quality with scale requires combining expert-created seed samples with automated methods [316]. For LLM-prompt benchmarks, one can include jailbreaking prompts to test worst-case performance. To ensure robust assessment quality, a combination of LLM judges, dynamic evaluation methods, and static program analysis techniques is recommended.
Benchmarks should reflect realistic scenarios. Real-world testing would be ideal, though this poses many challenges, including verification challenges and legal/ethical constraints. Efforts could focus on replicating real-world production systems with verification mechanisms to assess frontier AI capabilities. Complementary proxy benchmarks can also be valuable along with designing appropriate metrics—such as Cybench’s “first solve time” which compares agent performance to human expert solving times [329]. Finally, cybersecurity benchmarks require continuous updates due to its rapidly evolving threat landscape. Unlike relatively static domains like mathematics or history, cybersecurity threats transform quickly—agentic AI hacking, once considered science fiction, is now a significant concern. Therefore, even if tasks remain hidden and are not trained on, we must continuously invest in benchmarks relevant to emerging threats.
7.2 Enhance Cyber Defenses
Improve proactive testing. First, more efforts are needed in designing automated penetration testing benchmarks and agent-based tools. Regarding the agent, future efforts could focus on providing richer tool support, such as searching for new vulnerabilities in external databases or the internet and tools that help maintain access. With rich tools, we also need to design agents with stronger planning and more capabilities beyond the current ones, such as privilege escalation, compromising domain controllers, and lateral movements.
Second, for vulnerability detection, we recommend leveraging frontier AI to solve the key challenges of existing program analysis-based methods. For static analysis, promising directions include AI-based state pruning to alleviate state explosion and extracting static rules through model explanation to enhance knowledge bases. For fuzzing, one can explore efficient seed/mutator scheduling with RL and LLM-based grammar generation for domain-specific fuzzers. Third, we recommend fine-tuning customized models for better program state understanding and vulnerability reasoning. This requires high-quality training data with correct labels and more program artifacts and context, especially dynamic information, to teach the model program states [81]. We then need to design customized fine-tuning algorithms to gain testing-phase reasoning capabilities. Distillation of large reasoning models through supervised fine-tuning and reinforcement learning with proper reward functions are promising directions, which are effective in code generation and math problem-solving [168, 203, 112]. It is critical to scale AI-driven vulnerability detections beyond function levels to file and repository levels. Fine-tuning models with long context limits and building AI agents with program analysis tools (e.g., AST analysis and data/control flow analysis) could be promising directions.
Finally, we recommend constructing vulnerability detection benchmarks that avoid data leakages in the testing set and have balanced vulnerability distributions that consider the long tails. Having human experts work with AI can potentially strike a balance between reducing manual efforts and improving data quality.
Enhance attack detection. We first recommend constructing high-quality benchmarks that minimize the critical data and label issues, such as duplicated data, spurious correlations, biased samples, and wrong and missing labels. A recent research [18] provided a comprehensive summary of these issues. Given the challenges of collecting high-quality real-world attack data, collecting data from controlled environments [29] and creating synthetic data [302] could be efficient ways. It is necessary to construct hidden test sets that include unseen attacks. These hidden sets will also help continuously retrain AI-driven defenses (Once the currently hidden test sets are used for training, new test sets must be constructed and kept separate from the training data). Such practices have been applied in the natural language field: hiding test sets (e.g. SQuAD [228]), updating test sets with recent data (e.g. LiveBench [294]), or estimating and resolving training-test overlap [328].
Regarding the learning models and methods, although transformers address key challenges of other DNN models, it is still important to develop customized learning methods for security-specific challenges, e.g., attack drifting and extremely imbalanced classes, as well as improve the robustness of AI-driven detection methods through adversarial retraining and guardrails. When being used in critical scenarios with an extremely low error tolerance, combining AI-based attack detection with traditional detection methods is a promising direction to ensure reliability and robustness. For example, AI can categorize attack instances and generate insights from the data they’ve been trained on. Instead of manually analyzing every instance, security analysts can focus only on the representative cases from each category and extract static rules from AI.
Facilitate triage and remediation. We first recommend building multi-agent systems for automated triage and patching. For example, we can integrate differential fuzzing to validate the correctness and security of AI-generated patches. We can also integrate formal verifications in patching agents to provide formal functionality and security guarantees. It is important to enable iterative refinements for these agents based on feedback. Such agents can also be leveraged to improve the automation and efficiency of SOC and security analysts’ workflow (e.g., data filtering and management). Second, we recommend training specific AI models to reason about complex program artifacts (e.g., atypical constructs), dependencies (inter-procedure dependencies and implicit constraints), and vulnerabilities. The methods discussed above about improving proactive testing and trained models can be applied here. We can also leverage the agent’s feedback as the reward signal for model training, such as training models to generate secure patches based on the PoC testing. Third, we can train specialized small models for planning or tool calling and use them to replace LLMs used in classification and patching agents. Fourth, we recommend involving more AI components in remediation deployments, such as testing case generation, agentic-based configuration, and dependency analysis. We also recommend building CI/CD pipelines that integrate these strategies to systemically reshape defense deployment. Finally, we recommend constructing more security-specific benchmarks for triage and remediation that encompass complex code bases across multiple programming languages.
Improve the automation of defenses with provable guarantees. We believe frontier AI can significantly advance the scalability and automation of formal verification and call for increased efforts in this direction. First, we can use and fine-tune models to generate more complex proofs and constraints for theorem proving, as well as invariants for complex programs [314]. Second, we can use LLMs to generate necessary inputs for formal verification, such as program invariants and function summaries, where enabling verifiable generation with correctness guarantees is crucial. More broadly, secure code generation is a promising direction that can be used to turn unsafe latency code into safe counterparts. Early research has explored strategies like prefix learning [127] and constrained decoding [97]. We suggest focusing more on fundamental improvements in the models or agents, such as RL-based fine-tuning to improve models’ inherent capability of generating safe code. We can also design customized models and agents to transpile legacy programs into newer and memory-safe languages (e.g., C to Rust [87]). Third, AI can also improve constrained solvers. For example, SOTA solvers rely on somewhat random/rule-based mechanisms for backtracking and re-selecting verification paths to resolve conflicts. With frontier AI, solvers could learn more efficient strategies for making these critical decisions. Alongside building better AI-enhanced defenses, society must improve overall security posture and actively deploy these new defensive technologies.
7.3 Build Secure Hybrid Systems
We call for more effort in developing secure mechanisms for hybrid systems, whether through secure-by-design and secure-by-construction approaches or by creating specialized proactive and reactive defenses. This is essential to address the significant gap between the rapid emergence of hybrid systems and the limited exploration of their security. First, the security community has developed various protection mechanisms for traditional symbolic systems, such as strong authentication, access control, and secure protocol designs. However, with the rise of hybrid systems, there is currently no standardized secure design framework or principles tailored for such systems. As such, we recommend designing secure system design frameworks and principles for hybrid systems. For example, these new frameworks and standards could adapt existing security enforcement mechanisms (e.g., privilege isolation), as well as create new mechanisms that are customized for the marginal risks unique to hybrid systems (e.g., tool permission controls and specific sandboxing). For example, having a more fine-grained rigorous isolation mechanism for AI systems also helps with safety and security, e.g., separating system prompts from user prompts helps mitigate prompt injections [51].
Second, we recommend building real-time monitoring and alert mechanisms for hybrid systems. For example, we can extend existing AI attack detectors [9] to frontier AI, improve their efficiency, and continuously update them. In the long term, achieving provable security guarantees for hybrid systems is crucial. One possible direction is to follow a divide-and-conquer approach: decompose the verification of a hybrid system into sub-goals for its AI and traditional symbolic components, applying or extending specific techniques to address each. Overall, this remains an open challenge requiring greater attention and collaborative efforts from multiple computer science communities, including security, AI, programming languages, etc.
7.4 Recommendations for AI Development
Testing the cyberattack capabilities of new models before release and publishing detailed evaluation reports remains essential, as demonstrated by the GPT system card [140] and the Gemini dangerous capability evaluation [221]. As benchmarks evolve, AI developers should conduct more comprehensive testing beyond CTF competitions and collaborate with white-box hackers for real-world scenario testing. We can also explore implementing finer-grained privilege and access controls to ensure responsible use and avoid equivalence classes, enabling established defenders to access frontier AI more readily than attackers. However, implementation challenges remain, particularly in defining “established defenders.” While some arguments support open-source frontier AI to foster innovation, decisions about open-sourcing models and weights require careful risk assessment [153, 232].
Safety and security demand continued and increased attention. We recommend extending current comprehensive red-teaming practices more broadly across AI development, particularly for hybrid systems. For blue-teaming, we advocate a best-effort approach incorporating complementary strategies: model safety alignment, guardrails for models and agents, and system-level protections. Transparency remains a key limitation of frontier AI [232]. Considering the challenges in explaining complex model decisions, transparency in development processes is crucial for informed policymaking. Implementing varied transparency levels for different entities (public, trusted third parties, government) could balance information sharing risks and benefits, though this approach raises significant implementation challenges [232].
7.5 Mitigation of Human-Related Risks
To mitigate human-related attack risks, we first call for better user and developer education about AI attacks and security practices. Human awareness remains critical in cybersecurity, with research showing that developers’ security knowledge gaps and user unawareness significantly increase vulnerability and damage [20, 22]. Studies demonstrate AI can enhance educational efficiency across domains [50, 241, 226], suggesting the potential of leveraging frontier AI for security education. AI-driven personalized learning systems can deliver more efficient security training tailored to individual knowledge levels [264, 176]. Intelligent nudge mechanisms represent another effective approach [123, 163], where AI systems detect potential phishing or scams and provide timely alerts, enhancing real-time threat recognition while minimizing user fatigue.
However, education alone is insufficient, as users often remain the weakest link. As detailed in Sections 3.2 and 4.3, AI-powered attacks affect humans on a massive scale, while defensive measures lag behind rapidly evolving threats. We urgently need defensive research addressing frontier AI-enabled sophisticated attacks rather than focusing solely on traditional scenarios.
LLM-enhanced malicious bot activities, including misinformation campaigns and fraud schemes, exemplify this challenge. AI-based protective measures are essential against these advanced threats [94]. Such mechanisms can analyze behavioral patterns, communication styles, and interaction frequencies to distinguish legitimate users from sophisticated bots. Platforms can implement intelligent screening mechanisms to filter out malicious bots attempting social engineering or fraud on social media.
Additionally, frontier AIs can create “honeypot” systems that attract and isolate these bots [230], engaging them in resource-draining interactions while learning their behaviors to develop better countermeasures. Given the rapid advancement of human-targeting attacks, game-theoretical approaches analyzing interactions between offensive and defensive frontier AI can provide valuable insights into this evolving security landscape.
8 Open Research Challenges and Questions
As emphasized in Section 7, community efforts are essential for preparing for the arrival of frontier AI in cybersecurity. Here, we provide key research challenges and questions from two aspects: attack and marginal risk assessment; empirical defenses and secure-by-design.
Attack and Marginal risk assessment.
  1. 1.
    How will frontier AI impact each attack step, especially given its rapid capability evolvement? How different are the levels of AI capabilities required for different steps for better automation of attack steps?
  2. 2.
    How will the attack capabilities augmented by AI be translated into actual risks in the real world? For example, will the increase of individual attack steps always increase the overall risks in the real world?
  3. 3.
    How will the increased AI capabilities change the economics of attacks? How will this change the attack stages and viable attack strategies? What new attack methods, steps, and strategies may emerge due to the changed AI capabilities and economics of attacks?
  4. 4.
    How will frontier AI impact attackers with different levels of resources and capabilities?
  5. 5.
    What AI systems with which capabilities will trigger the explosive applications of AI in attacks, similar to how ChatGPT revolutionized the NLP domain? And when is this likely to happen?
  6. 6.
    How should we best assess marginal risks via benchmarks? To what extent will benchmarks translate to real-world risks? How do we close the gap and design benchmarks to better measure real-world risks on an ongoing basis?
  7. 7.
    How will AI affect attackers’ capabilities of evading current attack detection systems?
Frontier AI for cyber defense. Questions 1$\sim$4) in the attack side can also be applied to the defense side (e.g., how will frontier AI impact defenders with different levels of resources and capabilities?). Additionally, we present the following key questions.
  1. 1.
    Which security defenses will be the first to fail as AI-driven attacks grow in capabilities? How can we better prepare for these challenges? How can we improve the broken defenses with AI or other tools?
  2. 2.
    How will AI enhance vulnerability detection in the future? Will it identify similar bugs as traditional program analysis-based methods or will it find different ones? How to maximize the effectiveness of AI and traditional approaches when used together?
  3. 3.
    Given their statistical nature, will different AI-based systems find different bugs in different runs? If this is true, is there a possibility that attackers may leverage this property to launch attacks given that defenders may find different bugs from them even if they are using similar AI tools?
  4. 4.
    Will AI tools reach a saturation point where they can only identify vulnerabilities within a certain budget, making it increasingly difficult—and exceedingly costly—to discover new ones beyond that threshold?
  5. 5.
    Is it possible for frontier AI to generate secure program patches with formal security guarantees? How to automate the patching process to reduce the patch deployment cycle? Is it possible to develop secure-by-design methodologies for hybrid systems?
  6. 6.
    At what point will the equivalence classes be broken, given that defenders gain access to stronger AI before attackers? Will this make launching effective attacks more difficult? For example, if frontier AI can conduct comprehensive end-to-end penetration testing, does that mean that attackers cannot leverage the same AI to launch attacks?
  7. 7.
    How to enable proper open source strategies to balance innovation, IP protection, transparency, and security?
9 Conclusion
This paper presents an in-depth analysis of frontier AI’s impact on cybersecurity and provides recommendations to security researchers, practitioners, and AI developers. This work highlights a few future directions. First, we will continuously update our qualitative and quantitative analyses to maintain an up-to-date understanding of AI’s evolving role in cybersecurity. Second, as AI is increasingly used in the physical world, future works can conduct a similar analysis for cyber-physical systems and broader hardware systems.